OpenStudy (anonymous):

can anyone integrate this (1+x^(-2/3))^1/2 I just can't get it my substitution won't work

OpenStudy (anonymous):

this is difficult

OpenStudy (amistre64):

\[\int u^{1/2}du=\frac 23 u^{3/2}\] in essense

OpenStudy (anonymous):

Nope,,, it is easy, use the substitution.. t=x^(-1/3)...

OpenStudy (amistre64):

\[\frac{d}{dx}(1+x^{-2/3})=\frac{-2}{3}x^{-5/2}\]bummer

OpenStudy (amistre64):

alomst thought that out right lol

OpenStudy (anonymous):

(1 + x^{2/3})^{3/2}

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=integrate++%281%2Bx%5E%28-2%2F3%29%29%5E1%2F2 wolf gives some steps but it looks like a pain to me

OpenStudy (anonymous):

lucifur how did you get that? that's the right answer but what substitution did you do to get it?

OpenStudy (anonymous):

take t = x^2/3

OpenStudy (anonymous):

take t = x^2/3 dx = (3/2) t^-1/2 * dt